OpenStudy (crashonce):

A cone is found by bending a sector of angle 120 degrees and radius 6 cm. The radius of the base of the cone is?

OpenStudy (crashonce):

@mathmath333

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@Compassionate @Hero

OpenStudy (perl):

The circumference of the base of the cone will be the arc length of the original sector

OpenStudy (crashonce):

oh yea

OpenStudy (crashonce):

wats the radians formula again? i forgot

OpenStudy (perl):



OpenStudy (perl):

arclength = radius * theta radians s = r *theta

OpenStudy (perl):

so change 120 degrees to radians 120 deg * ( pi rad / 180 deg)

OpenStudy (perl):

120 deg = 2/3 Pi

OpenStudy (perl):

s = 6 * (2/3 * Pi ) = 4 Pi

OpenStudy (perl):

so C = 4Pi but C = 2Pi * R so can you solve for R

OpenStudy (perl):

since "The circumference of the base of the cone will be the arc length of the original sector" after you bend the sector to form a cone arclength of original sector = circumference of cone